Louisiana State University Health Sciences Center Administration & Finance
 

Thursday, June 21, 2007

Dental School :

  • The installation of the new ceiling grid and tile continued in the Clinic Building.
  • The test results from the remediation project are pending for the remaining floors and AHUs that have not yet been cleared.  We anticipate that additional results will be coming in shortly.
  • Boe-Tel is continuing to pull network cable and making connections in offices and clinic spaces as the enhancement to the network infrastructure in both the Clinic and Administration Building progress.
    • The construction of a new network/telephone closet on the 2nd floor of the Clinic Building is proceeding.
  • The demolition phase of the Central Sterilization project is well underway.
  • The repairs to the exterior site lighting and the electrical work necessary to power the parking lot controls are continuing.
    • Old subsurface wires have been pulled and the damaged light standards have been dismantled. 
  • The replacement doors and frames needed for the 1st floors of the two buildings have been delivered and the installation work has started.
  • A change order proposal to replace a number of HVAC inline dampers in the buildings has been received and approved.  The contractor is ordering the needed parts and materials.
  • A change order proposal to make repairs to the compressed air line running through the utility tunnel from the Power Plant to the Clinic Building has been received and approved.  Repairs should begin shortly.
  • A proposal to replace the 1st floor ceilings in the Clinic and Administration Buildings has been prepared and submitted. 
    • The proposal also addresses the creation of two temporary rooms on the 1st floor of the Clinic Building for use by Facility Services and Supply Chain Management when the School of Dentistry returns to the campus.
  • The architects have submitted a proposal for additional testing of the roofs on the two buildings.  This testing will assist in accurately determining the extent and nature of the storm related damage. 
    • Once the scope of the damage is aligned with FEMA, plans will be developed and bid to complete the permanent repairs.

Morgue Cooler:  We have initiated a project to replace the cooler in the morgue located in the MEB.


  • Bids for a new cooler were solicited and the contract was awarded.
  • The cooler is currently being manufactured. 
  • The installation of the new cooler is being planned to begin in mid-August.
Multi-Disciplinary Labs AV Equipment: Significant enhancements to and modernization of the audio/visual equipment are being made in the Multi-Disciplinary Labs (MDLs) located on the 4th floor of the MEB.
  • All of the new LCD monitors, high resolution projectors, and motorized screens have been installed in Labs #2, #3, #4, #5, and #6.
  • The main switch and the centralized control center are being wired; connecting all of the devices together.

    Student Parking Lots:

  •  The contractor is preparing both Student Lot #2 & #3 for the application of the asphalt.

 

Cafeteria Renovation:

  • The new flooring is being put down and the remaining new walls are being framed.
  • The new air handling unit has been delivered and installed.